# The basis problem for modular forms for the Weil representation

### v1

@@ 7 Applications
Waldspurger’s result on the scalar valued basis problem can be derived from Theorem 6.7 by showing that any newform of level $N$ is the 00-component of a suitable vector valued cusp form. For a positive integer $N$ let $\mathrm{S}_{k}^{\text{new}}(N)$ denote the space of scalar valued newforms of level $N$. Following [Wal] we let $\Theta(m,k,N,D)$ denote the space generated by scalar valued theta series of positive-definite even lattices of rank $m$, level $N$ and discriminant $D$ weighted with harmonic polynomials of degree $k-m/2$, i.e. the 00-components of elements in $\Theta_{m,k}(\mathcal{D})$, where $\mathcal{D}$ has level $N$ and $|\mathcal{D}|=D$.
@@
    For $m=0\bmod 8$ and $m\geq 8$ we have $\mathrm{S}_{k}^{\text{new}}(N)\subset\Theta(m,k,N,N^{2})$ (cf. [Wal, Theorem 1]).
@@
    For $m=4\bmod 8$ and $m\geq 12$ and for any prime $q\mid N$ we have $\mathrm{S}_{k}^{\text{new}}(N)\subset\Theta(m,k,N,N^{2}q^{2})$ (cf. [Wal, Theorem 2]).
@@
and consider the map $\Psi_{\mathcal{D}}:\mathrm{S}_{k}^{\text{new}}(N)\to\mathrm{S}_{k}(N)$ defined by $\Psi_{\mathcal{D}}(f)=\langle\mathcal{L}_{\mathcal{D}}(f),e^{0}\rangle$. In [SV, Theorem 1.1] it was shown that for certain discriminant forms, $\Psi_{\mathcal{D}}$ is a non-zero multiple of the identity map, say $c_{\mathcal{D}}\cdot\mathrm{id}$.
@@
Then $I\!I_{m,0}(\mathcal{D})$ is non-empty and satisfies the conditions of Theorem 6.7 as well as those of [SV, Theorem 1.1], so that for any $f\in\mathrm{S}_{k}^{\text{new}}(N)$ we find $\mathcal{L}_{\mathcal{D}}(f)\in\Theta_{m,k}(\mathcal{D})$ and hence, $f=\Psi_{\mathcal{D}}(c_{\mathcal{D}}^{-1}f)\in\Theta(m,k,N,N^{2})$.
@@
where $u$ is a non-square modulo $q$. Then $I\!I_{m,0}(\mathcal{D})$ is non-empty and again, the conditions of Theorem 6.7 and those of [SV, Theorem 1.1] are satisfied, which implies $\mathrm{S}_{k}^{\text{new}}(N)\subset\Theta(m,k,N,N^{2}q^{2})$. ∎
Theorem 3 in [Wal] on cuspforms with character can probably also be proved using Theorem 6.7, but we have not checked all the details.
@@
Now we want to apply the main result of the present work to show that the space of local obstructions for constructing Borcherds products generates the space of global obstructions. This extends Theorem 5.4 in [BF]. For details on Borcherds products see for example [Bo1], [Bo2], [Br] and [BF].
Let $L$ be an even lattice of signature $(l,2)$ with bilinear form (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) and quadratic form $\operatorname{q}(x)=\frac{1}{2}(x,x)$. Note that in [BF] lattices of signature $(2,l)$ were considered. Throughout we will assume that $l>8$ is even and that $L$ splits two hyperbolic planes $I\!I_{1,1}\oplus I\!I_{1,1}$. The complex manifold
@@
where $U$ ranges through all open neighbourhoods of $s$ and $U_{\mathrm{reg}}=U\cap(\Gamma^{+}\backslash\mathcal{H}_{l})$. For $\Gamma_{\infty}=P\cap\Gamma$, where $P\subset\operatorname{O}(L\otimes_{\mathbb{Z}}\mathbb{R})$ is the parabolic subgroup stabilizing $F\otimes\mathbb{R}$, we can consider the pullback of divisors on $\Gamma^{+}\backslash\mathcal{H}_{l}$ to divisors on $\Gamma_{\infty}^{+}\backslash U$. The pullback of $H(\beta,n)$ to $\mathrm{Pic}(X_{\Gamma},s)$ is denoted by $H_{F}(\beta,n)$ and will be called a local Heegner divisor. One can show that
$H_{F}(\beta,n)=\sum_{\begin{subarray}{c}\lambda\in(\beta+L)\cap F^{\bot}\\ \operatorname{q}(\lambda)=n\end{subarray}}\lambda^{\bot}$
(for details see [BF]). The following proposition was shown in [BF].
Proposition 7.2 (Proposition 5.1, [BF]).
@@
for all theta series $\theta_{M,P}$ with spherical polynomial $P\in\operatorname{H}_{l-2}^{2}$ and Fourier coefficients $a(\beta,n)$ ($\gamma\in M^{\prime}/M$ and $n\in\mathbb{Z}+\operatorname{q}(\gamma)$).
Let from now on $\Gamma=\ker(\operatorname{O}(L)\to\operatorname{O}(L^{\prime}/L))$ be the discriminant kernel of $L$. In [Bo1] Borcherds constructed lifts of certain vector valued modular forms of weight $21-l/$ 2 for $\mathrm{Mp}_{2}(\mathbb{Z})$ to meromorphic modular forms for the group $\Gamma^{+}$ (cf. Theorem 13.3 [Bo1]). These lifts have product expansions at 00-dimensional cusps and are therefore called automorphic products or Borcherds products. Their divisors are linear combinations of Heegner divisors. The following characterization can be found in [BF, Theorem 5.2] and goes back to Borcherds (cf. Theorem 3.1 [Bo2]).
@@
(with $c(\beta,n)\in\mathbb{Z}$ and $c(\beta,n)=c(-\beta,n)$) is the divisor of a Borcherds product for the group $\Gamma^{+}$ (as in [Bo1, Theorem 13.3]) if and only if for any cusp form $f\in\mathrm{S}_{1+l/2}(\mathcal{D})$ with Fourier coefficients $a(\beta,n)$ the equality
@@
The following result was suggested by Jan Bruinier and generalizes [BF, Theorem 5.4] to non-unimodular lattices.
@@
    $H$ is the divisor of a Borcherds product for the group $\Gamma^{+}$ as in [Bo1, Theorem 13.3].
@@
The modularity of a meromorphic modular form $\psi$ for the group $\Gamma^{+}$ immediately implies that the divisor $(\psi)$ attached to $\psi$ is trivial at generic boundary points. Therefore, we only need to prove (iii) implies (i). So assume that $H$ is trivial at generic boundary points. Note that, since $L$ splits two hyperbolic planes, the genus of $L$ contains only one class and the natural projection $\operatorname{O}(L)\rightarrow\operatorname{O}(L^{\prime}/L)$ is surjective (cf. [N, Theorem 1.14.2]). So by proposition 7.2
@@
As a corollary we find for lattices $L$ that satisfy the conditions of Theorem 7.5 that any meromorphic modular form for the discriminant kernel of $L$, whose divisor is a linear combination of Heegner divisors, is a Borcherds product. This was already proved in greater generality in [Br], however, using an entirely different argument, which says nothing about the local Picard groups.

### v2

@@ 7 Applications
Waldspurger’s result on the scalar valued basis problem can be derived from Theorem 6.7 by showing that any newform of level $N$ is the 00-component of a suitable vector valued cusp form. For a positive integer $N$ let $\mathrm{S}_{k}^{\text{new}}(N)$ denote the space of scalar valued newforms of level $N$. Following [33] we let $\Theta(m,k,N,D)$ denote the space generated by scalar valued theta series of positive-definite even lattices of rank $m$, level $N$ and discriminant $D$ weighted with harmonic polynomials of degree $k-m/2$, i.e. the 00-components of elements in $\Theta_{m,k}(\mathcal{D})$, where $\mathcal{D}$ has level $N$ and $|\mathcal{D}|=D$.
@@
    For $m=0\bmod 8$ and $m\geq 8$ we have $\mathrm{S}_{k}^{\text{new}}(N)\subset\Theta(m,k,N,N^{2})$ (cf. [33, Theorem 1]).
@@
    For $m=4\bmod 8$ and $m\geq 12$ and for any prime $q\mid N$ we have $\mathrm{S}_{k}^{\text{new}}(N)\subset\Theta(m,k,N,N^{2}q^{2})$ (cf. [33, Theorem 2]).
@@
and consider the map $\Psi_{\mathcal{D}}:\mathrm{S}_{k}^{\text{new}}(N)\to\mathrm{S}_{k}(N)$ defined by $\Psi_{\mathcal{D}}(f)=\langle\mathcal{L}_{\mathcal{D}}(f),e^{0}\rangle$. In [30, Theorem 1.1] it was shown that for certain discriminant forms, $\Psi_{\mathcal{D}}$ is a non-zero multiple of the identity map, say $c_{\mathcal{D}}\cdot\mathrm{id}$.
@@
Then $I\!I_{m,0}(\mathcal{D})$ is non-empty and satisfies the conditions of Theorem 6.7 as well as those of [30, Theorem 1.1], so that for any $f\in\mathrm{S}_{k}^{\text{new}}(N)$ we find $\mathcal{L}_{\mathcal{D}}(f)\in\Theta_{m,k}(\mathcal{D})$ and hence, $f=\Psi_{\mathcal{D}}(c_{\mathcal{D}}^{-1}f)\in\Theta(m,k,N,N^{2})$.
@@
where $u$ is a non-square modulo $q$. Then $I\!I_{m,0}(\mathcal{D})$ is non-empty and again, the conditions of Theorem 6.7 and those of [30, Theorem 1.1] are satisfied, which implies $\mathrm{S}_{k}^{\text{new}}(N)\subset\Theta(m,k,N,N^{2}q^{2})$. ∎
Theorem 3 in [33] on cuspforms with character can probably also be proved using Theorem 6.7, but we have not checked all the details.
@@
Now we want to apply the main result of the present work to show that the space of local obstructions for constructing Borcherds products generates the space of global obstructions. This extends Theorem 5.4 in [8]. For details on Borcherds products see for example [4], [5], [7] and [8].
Let $L$ be an even lattice of signature $(l,2)$ with bilinear form (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) and quadratic form $\operatorname{q}(x)=\frac{1}{2}(x,x)$. Note that in [8] lattices of signature $(2,l)$ were considered. Throughout we will assume that $l>8$ is even and that $L$ splits two hyperbolic planes $I\!I_{1,1}\oplus I\!I_{1,1}$. The complex manifold
@@
where $U$ ranges through all open neighbourhoods of $s$ and $\ΓU_{\mathrm{reg}}=U\cap(\Gamma\backslash\mathcal{H})$. For $\Gamma_{\infty}=P\cap\Gamma$, where $P\subset\operatorname{O}(L\otimes_{\mathbb{Z}}\mathbb{R})$ is the parabolic subgroup stabilizing $F\otimes\mathbb{R}$, we can consider the pullback of divisors on $\Γ\Gamma\backslash\mathcal{H}$ to divisors on $\Gamma_{\infty}\backslash U_{\epsilon}$ for certain neighbourhoods $U_{\epsilon}$ (cf. [8, Section 4]). The pullback of $H(\beta,l)$ is denoted by $H_{F}(\beta,l)$ and will be called a local Heegner divisor. One can show that
$H_{F}(\beta,l)=\sum_{\begin{subarray}{c}\lambda\in(\beta+L)\cap F^{\bot}\\ \operatorname{q}(\lambda)=l\end{subarray}}\lambda^{\bot}$
and $H_{F}(\beta,l)$ defines an element of $\mathrm{Pic}(X_{\Gamma},s)$ (for details see [8]). The following proposition was shown in [8].
Proposition 7.2 (Proposition 5.1, [8]).
@@
for all theta series $\theta_{M,P}$ with spherical polynomial $P\in\operatorname{H}_{l-2}^{2}$ and Fourier coefficients $a(\beta,n)$ ($\beta\in M^{\prime}/M$ and $n\in\mathbb{Z}+\operatorname{q}(\beta)$).
Let from now on $\Gamma=\ker(\operatorname{O}(L)\to\operatorname{O}(L^{\prime}/L))$ be the discriminant kernel of $L$. In [4] Borcherds constructed lifts of certain vector valued modular forms of weight $21-l/$ 2 for $\mathrm{Mp}_{2}(\mathbb{Z})$ to meromorphic modular forms for the group $\Gamma^{+}$ (cf. Theorem 13.3 [4]). These lifts have product expansions at 00-dimensional cusps and are therefore called automorphic products or Borcherds products. Their divisors are linear combinations of Heegner divisors. The following characterization can be found in [8, Theorem 5.2] and goes back to Borcherds (cf. Theorem 3.1 [5]).
@@
(with $c(\beta,n)\in\mathbb{Z}$ and $c(\beta,n)=c(-\beta,n)$) is the divisor of a Borcherds product for the group $\Gamma^{+}$ (as in [4, Theorem 13.3]) if and only if for any cusp form $f\in\mathrm{S}_{1+l/2}(\mathcal{D})$ with Fourier coefficients $a(\beta,n)$ the equality
@@
The following result was suggested by Jan Bruinier and generalizes [8, Theorem 5.4] to non-unimodular lattices.
@@
    $H$ is the divisor of a Borcherds product for the group $\Gamma^{+}$ as in [4, Theorem 13.3].
@@
The modularity of a meromorphic modular form $\psi$ for the group $\Gamma^{+}$ immediately implies that the divisor $(\psi)$ attached to $\psi$ is trivial at generic boundary points. Therefore, we only need to prove (iii) implies (i). So assume that $H$ is trivial at generic boundary points. Note that, since $L$ splits two hyperbolic planes, the genus of $L$ contains only one class and the natural projection $\operatorname{O}(L)\rightarrow\operatorname{O}(L^{\prime}/L)$ is surjective (cf. [25, Theorem 1.14.2]). So by proposition 7.2
@@
As a corollary we find for lattices $L$ that satisfy the conditions of Theorem 7.5 that any meromorphic modular form for the discriminant kernel of $L$, whose divisor is a linear combination of Heegner divisors, is a Borcherds product. This was already proved in greater generality in [7], however, using an entirely different argument, which says nothing about the local Picard groups.